OpenStudy (openstudyrocks5*):

The mean of four numbers is 30. If three of the numbers are 12, 50, 48, what is the fourth number?

OpenStudy (anonymous):

\[\frac{12+50+48+?}{4}=30\]

OpenStudy (openstudyrocks5*):

@iGreen

OpenStudy (igreen):

Follow @satellite73 's equation. \[\frac{12+50+48+x}{4}=30\] First we can add up 12 + 50 + 48, what do you get?

OpenStudy (openstudyrocks5*):

110 @iGreen

OpenStudy (igreen):

Yes, so we have: \(\dfrac{110+ x}{4} = 30\) Now we multiply 4 to both sides, what's 30 * 4?

OpenStudy (openstudyrocks5*):

120

OpenStudy (igreen):

Yes, so we have: \(110 + x = 120\) Now subtract 110 to both sides.

OpenStudy (openstudyrocks5*):

10?? @iGreen

OpenStudy (igreen):

Yep, you got it.

OpenStudy (igreen):

x = 10

OpenStudy (openstudyrocks5*):

Thx